what is the answer and how do you solve this problem -8=-2/3x.

Answer by RAY100(1637)

If problem is ,,,,-8 = -(2/3)x
.
multiply thru by -(3/2)
.
(-8)(-3/2) = (-2/3)(-3/2) x
.
+ 24/2 = +12 = x
.
.
check, (-2/3) (12) = -8,,,,,ok
.
.
if problem is,,,,-8 = -2/(3x),,,,multiply thru by -3x
.
(-3x)(-8) = (-3x)(-2/{3x})
.
24x = 2
.
x=1/12
.
.
check,,,,,-2/(3*1/12) = -2/ (1/4) = -8,,,,,ok
.